# Segment Anything in 3D Gaussians

### v2

@@ Abstract
3D Gaussian Splatting has emerged as an alternative 3D representation of Neural Radiance Fields (NeRFs), benefiting from its high-quality rendering results and real-time rendering speed. Considering the 3D Gaussian representation remains unparsed, it is necessary first to execute object segmentation within this domain. Subsequently, scene editing and collision detection can be performed, proving vital to a multitude of applications, such as virtual reality (VR), augmented reality (AR), game/movie production, etc. In this paper, we propose a novel approach to achieve object segmentation in 3D Gaussian via an interactive procedure without any training process and learned parameters. We refer to the proposed method as SA-GS, for Segment Anything in 3D Gaussians. Given a set of clicked points in a single input view, SA-GS can generalize SAM to achieve 3D consistent segmentation via the proposed multi-view mask generation and view-wise label assignment methods. We also propose a cross-view label-voting approach to assign labels from different views. In addition, in order to address the boundary roughness issue of segmented objects resulting from the non-negligible spatial sizes of 3D Gaussian located at the boundary, SA-GS incorporates the simple but effective Gaussian Decomposition scheme. Extensive experiments demonstrate that SA-GS achieves high-quality 3D segmentation results, which can also be easily applied for scene editing and collision detection tasks. Codes will be released soon.
@@ 1 Introduction
Figure 1: We propose an effective 3D object segmentation method in 3D Gaussian without any training process. The segmented mask contributes to the realization of various applications, including collision detection, translation, rotation and removal, etc.
@@
Alternatively, SA3D  [3] addresses the aforementioned challenges by enhancing the 2D foundational model (i.e. SAM  [13]) with 3D perception through the incorporation of a 3D representation model. SA3D demonstrates notable success by leveraging NeRF to generate 2D views and aligning 2D segmentation results with the 3D scene. However, the substantial training and rendering time costs associated with NeRF make it impractical for large-scale 3D scenes. Consequently, in this work, we focus on 3D Gaussian Splatting (3D-GS) [11] as a prospective method for modeling static 3D scenes. 3D-GS characterizes intricate scenes by employing numerous colored 3D Gaussians, rendering them into camera views through splatting-based rasterization. Through differentiable rendering and gradient-based optimization, the positions, sizes, rotations, colors, and opacities of these Gaussians can be finely tuned to accurately represent the 3D scene, availing for the comprehension of a 3D environment.
Benefiting from the properties of high-quality and real-time rendering in 3D-GS, it is impressive to use these explicit representations to capture the shape and appearance of complex objects. Specifically, based on the 3D Gaussian volume, we can effectively achieve scene perception via equipping the 2D foundation model as SA3D  [3] statement. Different from SA3D leveraging implicit NeRF as a 3D prior to connecting multiple 2D views, our method (SA-GS) applies effect and explicit 3D-GS representation to perform 3D object segmentation. Subsequently, scene editing and collision detection can be executed readily, as shown in Figure 1, which is essential to various applications, such as virtual reality(VR), augmented reality(AR), and games/movie production. Moreover, in contrast to the existing labor-intensive data collection perception frameworks that require point or multi-view images, 3D-GS offers a source of high-quality training data without the need for extensive manual efforts. This aspect holds significant promise for future research endeavors.
In this paper, we propose a novel approach to achieve object segmentation in 3D Gaussian via an interactive procedure without any training process and learned parameters. Specifically, by leveraging the 2D foundational model SAM  [13], SA-GS can generate a segmented mask from a single input view based on the given point prompts. Starting from the obtained mask, we automatically generate multi-view masks and achieve consistent 3D segmentation via the proposed view-wise label assignment. In addition, to address the boundary roughness issue of segmented objects resulting from the non-negligible spatial sizes of 3D Gaussian located at the boundary, SA-GS incorporates the simple but effective Gaussian Decomposition scheme. Moreover, our segmented 3D Gaussians can be seamlessly employed in collision detection and scene editing tasks, like removal, translation, rotation, etc. In summary, the contributions of this paper are as follows.
@@
  • We propose a simple yet effective method for object segmentation in 3D Gaussians without any training process and learnable parameters.
  • Efficient collision detection and scene editing, such as object removal, translation, and rotation within 3D scenes, can be performed readily based on the segmentation of SA-GS.
@@
  • We conduct extensive segmentation experiments on a considerable amount of 3D scenes, demonstrating the effectiveness of our proposed method.
@@ 2 Related Work
3D Gaussian Splatting [11] is a technique for scene reconstruction, which has been shown to be an alternative 3D representation as NeRF, benefiting from both its high-quality rendering results and real-time rendering speed. Recent research on this technique contains applying 3D Gaussian in dynamic scene and combining it with diffusion model to achieve 3D generation. Dynamic 3D Gaussians [17] extends Gaussian Splatting to dynamic scenes via tracking 3D objects modeled as a set of 3D Gaussians. DreamGaussian [29], GaussianDreamer [32] and GSGEN [5] combine 3D Gaussians with diffusion models to generate high-quality 3D assets. In this work, we for the first time lift 2D foundation model-SAM[13] to 3D Gaussians in order to achieve 3D object segmentation in this Gaussian space.
2.2 Segmentation in NeRFs
Neural Radiance Fields (NeRFs) are a popular way of representing 3D scenes implicitly with neural networks. Many researchers have explored how to segment objects in 3D using NeRFs, for various applications such as novel view synthesis, semantic segmentation, 3D inpainting, and language grounding. Some methods, such as Semantic-NeRF [33], NVOS [26], and SA3D [3], use different types of inputs to guide the segmentation, such as semantic labels, user scribbles, or 2D masks. Other methods, such as N3F [30], DFF [30], LERF [12], and ISRF [8], learn additional feature fields that are aligned with NeRFs, and use 2D visual features from pre-trained models or language embeddings to query the 3D features. These methods usually require modifying or retraining the original NeRF models or training other specific parameters to obtain the 3D segmentation. However, limited by the representation and rendering speed of NeRF, it is challenging to apply it to more practical applications, such as scene editing, collision analysis, etc. Benefiting from the purely explicit representation of 3D Gaussian, we can directly achieve 3D segmentation via an interactive way without any learned parameters and training process, and can apply for many other applications.
Figure 2: Pipeline of our proposed SA-GS. (a) Given a set of clicked points on the $1^{st}$ rendered view, we utilize SAM to generate masks for corresponding objects under every view automatically; (b) For every view, Gaussian Decomposition is first performed to address the issue of boundary roughness and then label propagation is implemented to assign binary labels to each 3D Gaussian; (c) Finally, with assigned 3D labels from all views, we adopt a simple yet effective voting strategy to determine the segmented Gaussians.
@@ 3 Method
In this section, we first present the preliminary in 3D Gaussian Splatting in Sec. 3.1 for clear understanding. We then define the problem and task in Sec. 3.2. Sec. 3.3, Sec. 3.4 and Sec. 3.5 elaborate on the technical details.
@@ 3.3 Segment 3D Gaussians with 2D Mask
3D Prompts for Multiview Masks Generation
By the definition in Sec. 3.2, users are given the first rendered view to specify the target object. However, only the first view is far from sufficient to segment the target object in 3D space. So here we first generate multiview masks to aid the 3D segmentation. With multiple masks from different views, a 3D object can be segmented by the intersection of the corresponding frustum of these masks.
@@
View-wise Label Assignment
With all the masks, we proceed to assign binary labels to each 3D Gaussian. In particular, we maintain a matrix $\mathbf{L}$, whose element $\mathbf{L}_{ij}$ is defined by
@@ 3.4 Gaussian Decomposition
Eq. 3 assign labels to 3D Gaussians only by the projected position of its center. However, 3D Gaussian has non-negligible spatial volume, so those Gaussians projected to the mask boundary usually have a part out of the boundary, greatly increasing the roughness of boundaries. A straightforward solution is directly removing the Gaussians across mask boundaries. However, such a solution greatly damages the 3D structures of the object.
To address this issue, we propose the Gaussian Decomposition to mitigate the boundary roughness while maintaining the 3D structures as complete as possible. Figure. 3 illustrates the basic idea of Gaussian decomposition. It has two basic steps to achieve the Gaussian decomposition: (1) For each 3D Gaussian, we obtain the corresponding 2D Gaussian by projection and mark it as a boundary Gaussian if one of its long-axis endpoints is outside of the 2D mask. Then, as shown in Figure. 3 we denote the two ends of the long axis of a 2D Gaussian as $A^{*}$ and $B^{*}$, and the intersection of $A^{*}B^{*}$ and mask boundary is $O^{*}$. $A$, $B$, and $O$ are the corresponding 3D points in the long axis of the 3D Gaussian. Assuming $A^{*}$ is in the mask while $B^{*}$ is out of the mask, we define
@@
Let $\mathbf{g}$ denote a 3D Gaussian across the boundary. Its scale in long axis and the 3D center are defined as $s$ and $\mathbf{\mu}$, respectively. We have
@@ 4 Experimets
Figure 4: Visualization of segmentation results in different scenes (LERF-figurines [12], LERF-dozer-nerfgun-waldo [12], SPIn-NeRF-Orchids [21]). We enlarge the boxed area on the right for a better visualization.
@@ 4.2 Quantitative results
Scenes Single View MVSeg SA3D Ours
IoU Acc IoU Acc IoU Acc IoU Acc
Orchids 79.4 96.0 92.7 98.8 83.6 96.9 85.4 97.5
Ferns 95.2 99.3 94.3 99.2 97.1 99.6 92.0 98.9
Room 73.4 96.5 95.6 99.4 88.2 98.3 86.5 98.1
Horns 85.3 97.1 92.8 98.7 94.5 99.0 91.1 98.4
Fortress 94.1 99.1 97.7 99.7 98.3 99.8 96.6 99.5
Fork 69.4 98.5 87.9 99.5 89.4 99.6 83.4 99.3
Pinecone 57.0 92.5 93.4 99.2 92.9 99.1 92.0 99.0
Truck 37.9 77.9 85.2 95.1 90.8 96.7 93.0 97.9
Lego 76.0 99.1 74.9 99.2 92.2 99.8 88.4 99.7
Mean 74.1 95.2 90.5 98.8 91.9 98.8 89.9 98.7
Table 1: Quantitative results on SPIn-NeRF dataset
@@
Figure 5: Visualization examples of scene editing after the object segmentation. We offer three scene editing examples: removal, translation, and rotation.
In the comparison, it is noteworthy that both the MVSeg [21] and the SA3D [3] require additional parameters and a computation-costly training process. By contrast, the “Single view” [3] refers to mapping the 2D masks to the 3D space based on the corresponding depth information, which does not need additional training process. In this sense, our method is the same as the “Single view” that does not incorporate any additional training or model parameters. Results show that our approach can achieve comparable segmentation quality to the learnable MVSeg and SA3D methods. In some 360 outdoor scenes, such as “Truck”, our approach even outperforms the SA3D and the MVSeg. When compared with the training-free method “Single view”, our approach achieves a significant promotion of +15.8% IoU and +3.5% Acc. These results demonstrate our approach is very efficient in obtaining high-quality segmentation masks.
@@
Next, we replace the above point prompts with text prompts corresponding to the objects. Given a text input, we prompt the Grounding DINO [16] to generate target bounding boxes, which serve as input prompts for the SAM to obtain 2D segmentation masks. Then, following the segmentation process in Section 3, we can transform the 2D masks into 3D Gaussian masks corresponding to the text input. Similar to the above evaluation process, we also calculate the IoU and Acc between 2D rendered masks and given ground-truth masks. Quantitative results are shown in Table 2. For both room and truck scenes, the IoU and Acc values can achieve a similar level as using clicked points as input prompts. These results demonstrate our method has the ability to combine multi-modal prompts as input.
@@ 4.3 Qualitative results
Point-Guided Segmentation:
We first conduct visualization experiments on 3D object segmentation guided by one or a set of point prompts clicked on the first-view rendered image. The results are shown in Figure 4. Our approach demonstrates stable performance in diverse scenes and target objects, including multiple instance segmentation (1st row), salient object segmentation (2nd row), and partial segmentation (3rd row). Even when the target objects are of extremely small scales (the first row), our approach can still produce high-quality segmentation.
@@
Text-Guided Segmentation:
Next, we further conduct 3D object segmentation experiments guided by text prompts. We follow the same process as in Table 2 experiments to conduct visualization experiments. In order to compare our method with SA3D under the same setting, we choose the same garden scene from Mip-NeRF 360 dataset [2], using three text prompts including “The table”, “The vase” and “The bonsai”. Results in Figure 6 show that our approach can achieve accurate segmentation results by simply providing object names, which demonstrates the potential of our method in combining with language models. Compared with the SA3D in the “The table” case, our approach can segment the complete object with desired table legs, while SA3D only gives the tabletop.
Scene Editing:
Scene editing is a basic application for 3D reconstructed scenes. However, it’s quite difficult to do this without being able to locate the specific objects. This task demonstrates the ability of our approach to help edit the 3D scenes. Specifically, after segmenting the objects in the 3D Gaussian space, we can manipulate the objects by removing, translating, and rotating them. Thanks to the simplicity of the explicit Gaussian representations, without bells and whistles, our approach obtains satisfactory scene editing results, as shown in Figure 5. The instances in 1st column with red bounding boxes are objects to be segmented. It can be seen that with objects segmented in 3D Gaussians, they can be translated and rotated in any direction in the scene. After the removal of segmented objects, original scenes can still keep intact.
Collision Detection:
Collision detection is an indispensable component in practical 3D applications, such as games, movies, and simulators. In this task, we demonstrate our approach can directly help in revealing the collision body of target objects in the 3D Gaussian space. We choose two scenes from SPIn-NeRF dataset. Following the process of our segmentation method, we can obtain the corresponding segmented 3D Gaussian points in the scene. To this end, we use the Quickhull [1] algorithm to build the collision mesh upon our segmented objects. The results in Figure 7 show that our segmented objects can successfully derive correct convex hulls for downstream applications.

### v3

@@ Abstract
3D Gaussian Splatting has emerged as an alternative 3D representation for novel view synthesis, benefiting from its high-quality rendering results and real-time rendering speed. However, the 3D Gaussians learned by 3D-GS have ambiguous structures without any geometry constraints. This inherent issue in 3D-GS leads to a rough boundary when segmenting individual objects. To remedy these problems, we propose SAGD, a conceptually simple yet effective boundary-enhanced segmentation pipeline for 3D-GS to improve segmentation accuracy while preserving segmentation speed. Specifically, we introduce a Gaussian Decomposition scheme, which ingeniously utilizes the special structure of 3D Gaussian, finds out, and then decomposes the boundary Gaussians. Moreover, to achieve fast interactive 3D segmentation, we introduce a novel training-free pipeline by lifting a 2D foundation model to 3D-GS. Extensive experiments demonstrate that our approach achieves high-quality 3D segmentation without rough boundary issues, which can be easily applied to other scene editing tasks. Our code is public available at https://github.com/XuHu0529/SAGS.
Keywords:
3D Gaussian Splatting 3D Segmentation Boundary Issues
@@ 1 Introduction
Figure 1: (a)The training of 3D-GS doesn’t consider the structure of objects, leading to the ambiguous geometry; (b)Direct segmentation without Gaussian Decomposition processing will result in rough boundary segmentation; (3) The recent SAGA also has incomplete segmentation caused by the same issue; (d) Our full pipeline considers this issue and can achieve better segmentation.
Recently, 3D Gaussian Splatting (3D-GS) [12] is emerging as a prospective method for modeling static 3D scenes. 3D-GS characterizes intricate scenes by employing numerous colored 3D Gaussians, rendering them into camera views through splatting-based rasterization. Through differentiable rendering and gradient-based optimization, the positions, sizes, rotations, colors, and opacities of these Gaussians can be finely tuned to accurately represent the 3D scene, availing for the comprehension of a 3D environment. Segmentation in 3D-GS is quite a natural pathway to scene understanding for its explicit representation. Recent works [3, 34, 37, 16] have been proposed to achieve segmentation in 3D-GS via lifting 2D foundation models in a learnable way. They share the same ideas to distill other features or identity encodings generated from 2D foundation models to 3D Gaussian fields.
However, the Gaussians learned by 3D-GS are ambiguous without any geometry constraint, leading to the issue of boundary roughness. A single Gaussian might correspond to multiple objects, complicating the task of accurately segmenting individual objects (shown in Fig. 1 (a)). As shown in Fig. 1 (b), direct segmentation will leave rough edges at the boundary. This is because these Gaussians across multiple objects will be left if no additional processing is performed. In addition, even if SAGA [3] uses filtering and growing post-processing for segmented 3D Gaussians, its feature matching method also has the limitation of incomplete boundary, as shown in Fig. 1 (c).
We present our boundary-enhanced segmentation method, an interactive training-free pipeline for efficient and effective segmentation in 3D-GS without rough boundary issues. By leveraging the 2D foundation model SAM [14], our method can generate a segmented mask from a single input view based on the given prompts. Starting from the obtained mask, our method automatically generates multi-view masks and achieves consistent 3D segmentation via the proposed assignment strategy. Specifically, to resolve the inherent boundary issues, we incorporate a simple but effective Gaussian Decomposition scheme, ingeniously utilizing the special structure of 3D Gaussian. Since the boundary roughness issue results from the non-negligible spatial sizes of 3D Gaussians located at the boundary, our key insight is to find and then decompose the original boundary Gaussians according to our proposed principles. Consequently, our method eliminates the inherent issues and achieves more complete segmentation quickly and efficiently, as shown in Fig. 1 (d).
In summary, the contributions of this paper are as follows:
@@
  • We propose a simple yet effective training-free pipeline for segmentation in 3D Gaussians without any learnable parameters;
  • We incorporate the Gaussian Decomposition module to mitigate the boundary roughness issues in 3D segmentation resulting from inherent 3D-GS geometry;
@@
  • Extensive segmentation experiments on a considerable amount of 3D scenes and editing applications demonstrate the effectiveness of our proposed method.
@@ 2 Related Work
3D Gaussian Splatting [12] is a technique for scene reconstruction, which has been shown to be an alternative 3D representation as NeRF, benefiting from both its high-quality rendering results and real-time rendering speed. Recent research on this technique contains applying 3D Gaussian in dynamic scene and combining it with diffusion model to achieve 3D generation. Dynamic 3D Gaussians [19] extends Gaussian Splatting to dynamic scenes via tracking 3D objects modeled as a set of 3D Gaussians. DreamGaussian [31], GaussianDreamer [35] and GSGEN [6] combine 3D Gaussians with diffusion models to generate high-quality 3D assets.
2.2 Segmentation in Radiance Fields
Neural Radiance Fields (NeRFs) are a popular way of representing 3D scenes implicitly with neural networks. Many researchers have explored how to segment objects in 3D using NeRFs, for various applications such as novel view synthesis, semantic segmentation, 3D inpainting, and language grounding. Some methods, such as Semantic-NeRF [36], NVOS [28], and SA3D [4], use different types of inputs to guide the segmentation, such as semantic labels, user scribbles, or 2D masks. Other methods, such as N3F [32], DFF [32], LERF [13], and ISRF [9], learn additional feature fields that are aligned with NeRFs, and use 2D visual features from pre-trained models or language embeddings to query the 3D features. These methods usually require modifying or retraining the original NeRF models or training other specific parameters to obtain the 3D segmentation. However, limited by the representation and rendering speed of NeRF, it is challenging to apply it to more practical applications, such as scene editing, collision analysis, etc. Inspired by the real-time 3D-GS, recent works [3, 34, 37, 16] have been proposed to achieve segmentation in 3D-GS in a learnable way. SAGA [3] and Featrue 3DGS [37] distill the knowledge embedded in the SAM decoder into the feature field of 3D-GS, and Gaussian-Grouping [34] supervises the Identity Encodings during the differentiable rendering by leveraging the 2D mask predictions by SAM. However, no research has been proposed to solve the rough boundary issues in segmentation resulting from the inherent properties of 3D Gaussians. In this work, we are the first to propose a novel Gaussian Decomposition scheme to solve this issue, incorporated into our effective training-free segmentation pipeline.
Figure 2: Pipeline of our proposed method. (a) Given a set of clicked points on the $1^{st}$ rendered view, we utilize SAM to generate masks for corresponding objects under every view automatically; (b) For every view, Gaussian Decomposition is first performed to address the issue of boundary roughness and then label propagation is implemented to assign binary labels to each 3D Gaussian; (c) Finally, with assigned 3D labels from all views, we adopt a simple yet effective voting strategy to determine the segmented Gaussians.
@@ 3 Method
In this section, we first present the preliminary in 3D Gaussian Splatting in Sec. 3.1 for clear understanding and then define the problem and task in Sec. 3.2. Our basic training-free pipeline consists of Sec. 3.3 and Sec. 3.5. The details of Gaussian Decomposition are described in Sec. 3.4. The pipeline of our method can be seen in Fig. 2.
@@ 3.3 Segment 3D Gaussians with 2D Mask
3D Prompts for Multiview Masks Generation. By the definition in Sec. 3.2, users are given the first rendered view to specify the target object. However, only the first view is far from sufficient to segment the target object in 3D space. So here we first generate multiview masks to aid the 3D segmentation. With multiple masks from different views, a 3D object can be segmented by the intersection of the corresponding frustum of these masks.
@@
View-wise Label Assignment. With all the masks, we proceed to assign binary labels to each 3D Gaussian. In particular, we maintain a matrix $\mathbf{L}$, whose element $\mathbf{L}_{ij}$ is defined by
@@ 3.4 Gaussian Decomposition
After obtaining the prompt points of each view, we could assign labels to 3D Gaussians by the projected position of its center. However, 3D Gaussian has non-negligible spatial volume, so those Gaussians projected to the mask boundary usually have a part out of the boundary, greatly increasing the roughness of boundaries. A straightforward solution is directly removing the Gaussians across mask boundaries. However, such a solution greatly damages the 3D structures of the object.
[…]
@@
[…]
@@ 4.2 Quantitative results
Table 1: Quantitative results on SPIn-NeRF dataset. “Singe view” denotes projecting the 2D segmentation result to 3D simply, thus we don’t consider its time cost. There is extra time cost (about 10mins) for SAGA to distill 2D knowledge to 3D-GS, while other methods don’t need this process.
Scenes Single View [4] MVSeg [23] SA3D [4] SAGA [3] Ours
IoU Acc IoU Acc IoU Acc IoU Acc IoU Acc
Orchids 79.4 96.0 92.7 98.8 83.6 96.9 - - 85.4 97.5
Ferns 95.2 99.3 94.3 99.2 97.1 99.6 - - 92.0 98.9
Room 73.4 96.5 95.6 99.4 88.2 98.3 - - 86.5 98.1
Horns 85.3 97.1 92.8 98.7 94.5 99.0 - - 91.1 98.4
Fortress 94.1 99.1 97.7 99.7 98.3 99.8 - - 96.6 99.5
Fork 69.4 98.5 87.9 99.5 89.4 99.6 - - 83.4 99.3
Pinecone 57.0 92.5 93.4 99.2 92.9 99.1 - - 92.0 99.0
Truck 37.9 77.9 85.2 95.1 90.8 96.7 - - 93.0 97.9
Lego 76.0 99.1 74.9 99.2 92.2 99.8 - - 88.4 99.7
Mean 74.1 95.2 90.5 98.8 91.9 98.8 88.0 98.5 89.9 98.7
Time cost - 3-6mins 2-10mins 0.08s-0.9s within 15s
(/object) (+10mins)
@@
In the comparison, it is noteworthy that both the MVSeg [23] and the SA3D [4] require additional parameters and a computation-costly training process. While SAGA [3] shows nearly one-thousandth of the time compared with them, it needs extra training time (10 minutes per scene) to distill the knowledge embedded in the SAM decoder into the feature field. By contrast, the “Single view” [4] refers to mapping the 2D masks to the 3D space based on the corresponding depth information, which does not need additional training process. In this sense, our method is the same as the “Single view” that does not incorporate any additional training or model parameters. Results show that our approach can achieve comparable segmentation quality to the MVSeg and SA3D methods and better results than SAGA. The time cost of ours is much less than MVSeg and SA3D. Considering the training time of SAGA, the average segmentation cost of a single object is basically the same as our method. In some 360 outdoor scenes, such as “Truck”, our approach even outperforms the SA3D and the MVSeg. When compared with the training-free method “Single view”, our approach achieves a significant promotion of +15.8% IoU and +3.5% Acc. These results demonstrate our approach is very efficient in obtaining high-quality segmentation masks.
@@
Figure 4: Qualitative results compared with SA3D [4] and SAGA [3] in different scenes (LERF-figurines [13], SPIn-NeRF-Orchids [23], LERF-dozer-nerfgun-waldo [13]). We enlarge the boxed area on the right for a better visualization.
Next, we replace the above point prompts with text prompts corresponding to the objects. Given a text input, we prompt the Grounding DINO [18] to generate target bounding boxes, which serve as input prompts for the SAM to obtain 2D segmentation masks. Then, following the segmentation process in Section 3, we can transform the 2D masks into 3D Gaussian masks corresponding to the text input. Similar to the above evaluation process, we also calculate the IoU and Acc between 2D rendered masks and given ground-truth masks. Quantitative results are shown in Table 2. For both room and truck scenes, the IoU and Acc values can achieve a similar level as using clicked points as input prompts. These results demonstrate our method can combine multi-modal prompts as input.
@@ 4.3 Qualitative results
Point-Guided Segmentation: We first conduct visualization experiments on 3D object segmentation guided by one or a set of point prompts clicked on the first-view rendered image. The results are shown in Fig. 4. The first row shows the segmentation results of SA3D, SAGA and ours on LERF-figurines scene. The second row compares with SAGA, which distills the knowledge embedded in the SAM decoder into the feature field. Though the boundary issues in 3D-GS can be alleviated, incomplete segmentation occurs even after post-processing. In contrast, ours can solve the boundary problems while achieving more complete segmentation. More segmentation results are shown in the third row.
Text-Guided Segmentation: We further conduct 3D object segmentation experiments guided by text prompts. We follow the same process as in Table 2 experiments to conduct visualization experiments. In order to compare our method with SA3D under the same setting, we choose the same garden scene from Mip-NeRF 360 dataset [2], using three text prompts including “The table”, “The vase” and “The bonsai”. Results in Fig. 5 show that our approach can achieve accurate segmentation results by simply providing object names, demonstrating our method’s potential in combining with language models. Compared with the SA3D in the “The table” case, our approach can segment the complete object with desired table legs, while SA3D only gives the tabletop.
Comparison between 2D and 3D Segmentation: Our method improves the segmentation quality of SAM. Due to the sensitivity of SAM to the scene viewpoint, the 2D segmentation results might be incomplete under some views, but our 3D segmentation can still guarantee complete segmentation results via a multi-view ensemble (Fig. 6).
@@
Scene Editing: Scene editing is a basic application for 3D reconstructed scenes. However, it’s quite difficult to do this without being able to locate the specific objects. This task demonstrates the ability of our approach to help edit the 3D scenes. Specifically, after segmenting the objects in the 3D Gaussian space, we can manipulate the objects by removing, translating, and rotating them. Thanks to the simplicity of the explicit Gaussian representations, without bells and whistles, our approach obtains satisfactory scene editing results, as shown in Fig. 7. The instances in 1st column with red bounding boxes are objects to be segmented. It can be seen that with objects segmented in 3D Gaussians, they can be translated and rotated in any direction in the scene. After the removal of segmented objects, original scenes can still keep intact.
Figure 6: Comparison between 2D and 3D Segmentation.
Collision Detection: Collision detection is an indispensable component in practical 3D applications, such as games, movies, and simulators. In this task, we demonstrate our approach can directly help in revealing the collision body of target objects in the 3D Gaussian space. We choose two scenes from SPIn-NeRF dataset. Following the process of our segmentation method, we can obtain the corresponding segmented 3D Gaussian points in the scene. To this end, we use the Quickhull [1] algorithm to build the collision mesh upon our segmented objects. The results in the last column of Fig. 7 show that our segmented objects can successfully derive correct convex hulls for downstream applications.
Figure 7: Visualization examples of scene editing after the object segmentation. We offer three scene editing examples: removal, translation, and rotation. We further provide collision mesh computed after segmentation, as shown in the last column.
4.4 Ablation Study
Gaussian Decomposition:
Fig. 8 shows the visualization results for comparison. We select two representative scenes in which the junction/contact exists large-scale Gaussian points (can be seen in the 2nd column). This issue is alleviated after utilizing the proposed Gaussian Decomposition strategy. Though this idea is simple, the improvement is obvious.